OpenStudy (anonymous):

Marissa has $325 to spend on the office party. She spends 3/5 of the money on food and spends 1/4 on decorations. How much money does she have left for invitations?

OpenStudy (anonymous):

so here is what you do you started with a total of 325 \[325 \times \frac{ 3 }{ 5} \times \frac{ 1 }{ 4 }\] I thank thats how you do it....